OpenStudy (anonymous):

I'm stomped, can anyone help me with this problem? Emily rows six miles downstream in 1 hour and her friend Ashley, rowing 1 mile per hour faster, completes the return trip in 2 hours. Find the speed of the current (c) and each girl's rowing speed. If Emily and Ashley were rowing separately, who would complete their trip first and by how long? Round to the nearest hundredth, if necessary.

OpenStudy (gorv):

let speed of current=x rowing speed of emily= y rowing speed of Ashley =y+1 time taken for downstream by emly = 1hour=distane * speed=6 *(x+y) as current will inc speed during downstream 6x+6y=1.....................equ 1 ashley 2= 6*(y+1-x).....in return current will oppose the speed and dec it 1=3y+3-3x 3x-3y=3-1=2 3x-3y=2......................equ 2 multiplying equ 2 by 2 and subtract it from equ1 6x+6y=1 -6x+6y=-4 12y=1-4=3 y=3/12 y=1/4=0.25miles / hour

OpenStudy (anonymous):

Ok, So Ashley would complete the trip first right? Thx

OpenStudy (anonymous):

A Mathematica calculated solution is attached.